# Time-Dependent VAE for Building Latent Factor from Visual Neural Activity with Complex Dynamics

### v1

@@ 4.2 Experiments on Synthetic Data
Figure 2: Results of synthetic experiments. A. The true latent variables. Different colors correspond to different labels $u$, gradient from 00 to $2\pi$. B. The regressed latent variables of our model. C-G. The results of alternative models. H. The reconstruction scores of all models. Each black dot represents an individual run, and the red dot represents the average score of ten runs.
We first validate TiDe-SwapVAE on the task of reconstructing the ground truth of synthetic latent variables. The procedure for generating synthetic data follows some previous work [17, 36].
4.2.1 Synthetic Dataset and Experimental Setup
First, we generate labels $u_{i}$ from four uniform distributions on $[\frac{2i\times\pi}{4},\frac{(2i+1)\times\pi}{4}],i\in\{0,1,2,3\}$, in preparation for building four clusters. Second, for each cluster, we sample 2-dimensional latent variables $\mathbf{z}$ from independent Gaussian distribution with $(5\sin{u_{i}},5\cos{u_{i}})$ as mean and $(0.6-0.5|\cos{u_{i}}|,0.5|\cos{u_{i}}|)$ as variance. Third, we feed sampled latent variables into a RealNVP network [8] to form the firing rate of 100-dimensional observations and generate the synthetic neural activity from the Poisson distribution. Each cluster contains 4,000 samples (Figure 2A).
After generation, all samples are split to 80% for training and 20% for test. All models are set up with 32-dimensional latent variables and trained for 1,000 epochs with an optimizer of Adam and a learning rate of 0.0005. We then apply linear regression to map the latent variables of models to the true latent and report $R^{2}$ of the linear regression as the reconstruction score. Specially, since the synthetic data does not involve temporal dependence, the length of input sequences is set to 1 even for models that can handle sequential data.
4.2.2 Reconstruction Results
As shown in Figure 2, TiDe-SwapVAE reliably separates the different clusters as well as recovers the structure of true latent variables to form clear arcs. In contrast, some of the alternative models fail to construct similar structures although they separate clusters (Swap-VAE and CEBRA), and others even have difficulty distinguishing four clusters ($\beta$-VAE, LFADS, and pi-VAE111pi-VAE incorporates the label prior during training, but the inferenced latent variables are built without the label prior at the evaluation stage. This way is used in all subsequent experiments.). Quantitatively, the results of reconstruction scores also suggest that our model outperforms all alternative models (Figure 2H).
@@ 4.3 Experiments on Mouse Visual Cortex Data
We aim to investigate whether TiDe-SwapVAE can build meaningful latent representations to reveal the intrinsic correlation between neural activity and visual stimulation. Therefore, we utilize a subset of the Allen Brain Observatory Visual Coding dataset [27] for evaluation, which has been used in a variety of work, such as constructing brain-like networks [26], modeling functional mechanisms [1, 7], and decoding neural representations [25]. This dataset is collected by Neuropixel probes from 6 mouse visual cortical regions simultaneously, including VISp, VISl, VISrl, VISal, VISpm, and VISam. Notably, the neural activity was recorded while mice were passively viewing visual stimuli (natural scenes and natural movie) without any task-driven behavior.
The dataset contains 32 sessions, each for one mouse. Since the class of neurons responsive to natural scenes or movies is found in six visual regions, in this work we choose to analyze the neural activity of four mice that have as many neurons as possible (about 300, see Appendix C for exact numbers) with them evenly distributed across all regions.
Table 1: The decoding scores (%) for natural scenes classification on the mouse visual neural dataset. The standard error is computed for 10 runs with different random initialization.
Models Mouse 1 Mouse 2 Mouse 3 Mouse 4
PCA (baseline) 20.0 43.2 43.6 20.0
$\beta$-VAE 50.0$\pm$2.8 44.8$\pm$1.1 39.2$\pm$2.0 46.0$\pm$3.0
LFADS 43.6$\pm$4.4 48.4$\pm$2.9 48.0$\pm$2.0 47.6$\pm$1.5
pi-VAE 71.6$\pm$3.7 66.0$\pm$2.1 76.0$\pm$1.5 28.8$\pm$2.8
Swap-VAE 70.4$\pm$3.9 66.0$\pm$1.5 64.0$\pm$2.6 53.2$\pm$2.8
CEBRA 38.8$\pm$1.3 46.8$\pm$2.6 47.2$\pm$2.0 43.2$\pm$2.6
TiDe-SwapVAE 80.8$\pm$2.1 70.8$\pm$1.6 63.6$\pm$2.8 56.4$\pm$2.6
4.3.1 Visual Dataset under Static Natural Scenes Stimuli and Experimental Setup
@@
First, we focus on neural activity in response to natural scenes. There are 118 natural scenes presented in random order, each for 250ms and 50 trials. We select five scenes that elicit the strongest average responses in four mice for experiments. The neural activity in the form of spike counts is binned into 10ms windows so that each trial contains 25 time points. We randomly split each dataset (one mouse) across all trials into 80% for training, 10% for validation, and 10% for test, and ensure that each subset has an equal number of trials per scene.
In this experiment, we set all models to have 128-dimensional latent space and train them for 200 epochs. The optimizer is also set to Adam with a learning rate of 0.0001. At the training stage, for the three models dealing with sequence data (LFADS, CEBRA and TiDe-SwapVAE), each sample is sequential neural activity from 5 time points. For self-supervised learning, the offset of positive samples from target samples is within ±$\pm{3}$ time points. At the inference stage, we consider the temporal dependence and the assumption that the latent variables of TiDe-SwapVAE are generated by an n-order Markov chain. Consequently, for a target time point, we form a sequence data including its antecedent n time points and its own to compute the latent variables (here n=4). For LFADS, we apply the same approach. For CEBRA, following the original approach [25], we take the surrounding points centered on the target point to form a sequence of the same length as during the training stage and compute its latent variables. For other models ($\beta$-VAE, pi-VAE and Swap-VAE), each input sample is neural activity of an independent time point, both at the training and inference stages.
@@
To quantify the performance on decoding natural scenes, we first obtain the latent variables of the last 20 time points (50ms-250ms) in each trial, since there is a response latency in the mouse visual cortex for static stimuli [27]. Then, these latent variables are concatenated into a vector as latent representations of neural activity for that trial. We use the KNN algorithm, a nonparametric supervised learning method, to classify the latent representations of each trial, i.e., to decode the corresponding natural scenes. We search the number of neighbors in odd numbers from 1 to 20 and use the Euclidean distance metric. We fit the KNN using the training set and choose the best number of neighbors on the validation set with classification accuracy as the metric. The accuracy on the test set is reported as the decoding score.
Figure 3: Results on mouse visual cortex dataset under static natural scenes. All dimension reduction is done by tSNE. A. The 2-dimensional embedding of the latent representations of our model for each trial. Each color corresponds to a natural scene. Transparent dots denote trials from the training set. Hollow dots denotes trials from the test set. B. The 2-dimensional embedding of the latent trajectories of our model over time, averaged over all trials for each scene. The asterisk is the starting point. C. The decoding scores of the full, content and style latent spaces of our model (mean $\pm$ standard error). D-G. The same visualization as A and B for alternative models.
4.3.2 Results of Decoding Natural Scenes and Analysis of Latent Representations
As shown in Table 1, TiDe-SwapVAE achieves the highest decoding scores on three mouse datasets with a noticeable improvement over other models. We observe some meaningful phenomena through comparisons. The fact that pi-VAE introduces class labels of natural scenes in the training stage gives it high decoding scores, but not stably (Mouse 4). The performance of Swap-VAE is consistently good, which suggests that the separation of two types of latent spaces and the swapping operation are indeed effective. However, the other two models (LFADS and CEBRA) that take sequential data as input instead perform poorly on the downstream task, suggesting that their approaches (bidirectional RNNs and temporal filters) of extracting temporal neural features are inappropriate in this case. In contrast, the temporal stepwise computation of our model along the chronological order shows more reliability in exploiting the time dependence. We further evaluate the decoding scores of the content and style latent spaces of our model (Figure 3C) on Mouse 1. We find that the content space yields higher scores than the full space, while the style space shows poor discrimination. The results provide valid evidence for the disentanglement of two spaces of our model.
In addition to the quantitative analysis, we visualize the latent representations by embedding them in two dimensions using tSNE (Figure 3, we focus on Mouse 1 which yields the highest decoding scores. Results of the other three mice are shown in Appendix D). On the one hand, we apply dimension reduction to the representations of each trial (including all time points from 50ms to 250ms), to show the trial-to-trial performance on decoding natural scenes. On the other hand, we reduce the dimensions of a single time point and take the average across trials, to show the latent trajectories over time. For the results of TiDe-SwapVAE, we first observe that most trials are well separated for different classes, especially those from the test set. The latent trajectories capture a similar clear temporal structure with weak class information. For pi-VAE and Swap-VAE, while they are able to discriminate scene classes of the trials, the latter part of latent trajectories show varying degrees of entanglement. As for LFADS and CEBRA, their ability to encode temporal features of sequential neural activity results in explicit temporal evolution, but the latent representations of different classes are largely intermingled. These results suggest that our model effectively distinguishes between category information as well as explicitly captures temporal information from neural dynamics.
This experiment demonstrates the advantages of our model in decoding static natural scene stimuli and extracting fine neural dynamics. However, we also see that the decoding performance for different mice exhibits large differences and the latent representations of some trials are not able to unravel the category entanglement. This may be due to the fact that in the case of passive viewing, the neural activity of different mice and even of different trials from the same individual show huge variability.
4.3.3 Visual Dataset under Natural Movie Stimuli and Experimental Setup
We use neural activity in response to 30s movie stimuli to build the dataset. There are 10 trials and the frame rate of the movie is 30Hz. We bin the spike counts with a sampling frequency of 120Hz and align them with the movie timestamps, resulting in 4 time points for each frame. Similar to the neural dataset under natural scenes, we split the dataset into training, validation and test set.
In this experiment, all models are set up with 128-dimensional latent variables. The number of epochs, the optimizer and the learning rate are the same as for the dataset under natural scenes. For our model, at the training stage, we use neural activity from 4 time points as one sample (also for LFADS and CEBRA) and set the maximum absolute offset to 2. At the inference stage, following the approach in Section 4.3.1, the latent variables of a target time point are generated based on antecedent n time points (n=3) and its own. The preprocessing of input samples for the other models is also consistent with Section 4.3.1.
To quantify the performance on decoding natural movie frames, we compute the latent variables of 4 time points within each frame and take the average as latent representations for that frame. KNN is applied to predict the movie frames corresponding to the latent representations (900 frames in total, i.e., 900 classes). Following [25], we take the accuracy measured by considering the error between a predicted frame and the true frame within 1s (default) as a correct prediction. We similarly use the validation set to find the best number of neighbors and report the accuracy of the test set as the decoding score.
4.3.4 Results of Decoding Natural Movie and Analysis of Latent Representations
We find that TiDe-SwapVAE again performs better than alternative models on all four mouse datasets (Table 2). Especially, our model achieves significantly higher decoding scores than LFADS, which is also a sequential VAE. Furthermore, we analyze the decoding scores under different sizes of time window constraints (maximum difference between predicted and real frames) on Mouse 2. The results of Figure 4D show that our model consistently outperforms CEBRA and Swap-VAE across a broad range of window sizes, while CEBRA’s performance degrades faster at smaller size constraints.
We similarly reduce the dimensions of the latent representations of each movie frame for visualization. We set all frames within 1s as a group and show the trajectories of latent representations evolving over time for the middle 10s of the movie (Figure 4A-C, see Appendix D for the results of other parts of the movie). All models present well temporal structure along movie frames. Compared to the two other models, the representations of TiDe-SwapVAE show less overlap and entanglement between different groups.
Our model constructs meaningful latent representations related to the content and temporal structure of movie stimuli at large time scales. At small time scales, although a drop in decoding performance occurs, our model is still better than alternatives. Similar features of adjacent frames may be a factor that makes the latter problem challenging, which may require further exploration.
@@ 4.3.4 Results of Decoding Natural Movie and Analysis of Latent Representations
Table 2: The decoding scores (%, in 1s window) for natural movie frame classification on the mouse neural dataset. The standard error is computed for 10 runs with different random initialization.
Models Mouse 1 Mouse 2 Mouse 3 Mouse 4
PCA (baseline) 8.44 28.77 21.56 11.69
$\beta$-VAE 7.32$\pm$0.27 14.93$\pm$0.34 17.56$\pm$0.62 9.44$\pm$0.19
LFADS 9.29$\pm$0.32 26.36$\pm$2.18 25.07$\pm$1.58 11.92$\pm$1.10
pi-VAE 10.19$\pm$0.27 44.83$\pm$0.71 35.84$\pm$0.45 20.22$\pm$0.34
Swap-VAE 10.88$\pm$0.27 43.71$\pm$1.05 35.36$\pm$0.73 19.24$\pm$0.47
CEBRA 9.91$\pm$0.36 44.62$\pm$1.00 39.48$\pm$0.40 21.41$\pm$0.38
TiDe-SwapVAE 12.07$\pm$0.38 49.01$\pm$1.10 45.21$\pm$0.41 21.52$\pm$0.47
@@
Figure 4: Results on mouse visual cortex dataset under natural movie. All dimension reduction is done by tSNE. A-C. The 2-dimensional embedding of the latent representations of our model for each frame. Each color corresponds to all frames within 1s. Each small dots denote one frame. Large dots denote the average among all frames in 1s. The red dashed line connects all averages. D. The decoding scores (%) for natural movie frame classification across different time window constraints. Error bars indicate the standard error for 10 runs.

### v2

@@ 4.2 Experiments on Synthetic Data
Figure 2: Results on the synthetic non-temporal dataset. A. The true latent variables. B. The regressed latent variables of our model. C. The results of an alternative model (see Appendix F for other alternatives). D. The reconstruction scores of all models. Each black dot represents an individual run, and the red dot represents the average score of ten runs.
We first validate TiDeSPL-VAE on the task of reconstructing synthetic latent variables. We generate two synthetic datasets for considering different properties of visual neural activity. One is a non-temporal dataset generated from several sets of labels, resembling categories of visual stimuli. The other is a temporal dataset constructed by the Lorenz system to test for time dependence. The generating procedure for both synthetic datasets follows some previous work [45, 26, 16, 38]. A detailed description of the datasets and the model training implementation is presented in Appendix C. After training, we apply linear regression to map latent variables of models to the ground truth on the test set and report $R^{2}$ of the linear regression as the reconstruction score.
Results on the non-temporal dataset
As shown in Figure 2, TiDeSPL-VAE reliably separates the different clusters as well as recovers the structure of true latent variables to form clear arcs. In contrast, some of the alternative models fail to construct similar structures although they separate clusters (Swap-VAE; CEBRA, Figure 6D of Appendix), and others are even difficult to split four clusters ($\beta$-VAE, LFADS, and pi-VAE111pi-VAE incorporates the label prior during training, but inferred latent variables are built without the label prior at the evaluation stage. This way is used in all subsequent experiments.; Figure 6A-C of Appendix). Quantitatively, the reconstruction scores also suggest that our model outperforms all alternative models (Figure 2D).
Results on the temporal dataset
In the first row of Table 1, TiDeSPL-VAE performs significantly better than those models that process sequential data at each time point independently, and moderately better than LFADS that uses bidirectional RNNs to handle temporal data. These results demonstrate the superiority of our model in dealing with time-dependent data. Moreover, when we shuffle the time dimension for each trial data on the original dataset to obtain a dataset without time dependence (the second row of Table 1), the performance of our model and LFADS shows a drastic degradation. However, the other models are less affected, with only Swap-VAE and CEBRA suffering a degradation due to the use of time-jittered positive samples. This phenomenon further supports the above conclusions.
Table 1: The reconstruction scores of all models on the synthetic temporal dataset. The standard error is computed based on 5 runs with different random initialization.
$\beta$-VAE LFADS pi-VAE Swap-VAE CEBRA TiDeSPL-VAE
Original 0.036$\pm$0.027 0.573$\pm$0.043 0.167$\pm$0.012 0.193$\pm$0.004 0.242$\pm$0.001 0.629$\pm$0.016
Shuffled 0.025$\pm$0.022 0.020$\pm$0.006 0.209$\pm$0.007 0.146$\pm$0.008 0.051$\pm$0.005 0.038$\pm$0.007
@@ 4.3 Experiments on Mouse Visual Cortex Data
We utilize a subset of the Allen Brain Observatory Visual Coding dataset [36] for evaluation, which has been used in a variety of work, such as constructing brain-like networks [35], modeling functional mechanisms [3, 9], and decoding neural representations [34]. This dataset is collected by Neuropixel probes from 6 mouse visual cortical regions simultaneously, including VISp, VISl, VISrl, VISal, VISpm, and VISam. Notably, the neural activity was recorded while mice passively viewed naturalistic visual stimuli without any task-driven behavior.
@@
The dataset contains 32 sessions, each for one mouse. Since the class of neurons responsive to natural visual stimuli is found in six visual regions, in this work we choose to analyze the neural activity of five mice that have as many neurons as possible (about 300, see Appendix D for exact numbers) with them evenly distributed across all regions (the coefficient of variation for the number of neurons across six brain regions is below 0.5). We focus on neural activity in response to natural scenes and natural movies. As for natural scenes, there are 118 images presented in random order, each for 250ms and 50 trials. We select five scenes that elicit the strongest average responses for experiments. The neural activity in the form of spike counts is binned into 10ms windows so that each trial contains 25 time points. As for the natural movie, it is 30s long with a frame rate of 30Hz, presented for 10 trials. We bin the spike counts with a sampling frequency of 120Hz and align them with the movie timestamps, resulting in 4 time points for each frame. For both datasets, we randomly split each across all trials into 80% for training, 10% for validation, and 10% for test.
4.3.1 Experiments for Neural Activity under Static Natural Scene Stimuli
Experiment setup and evaluation
@@ 4.3.1 Experiments for Neural Activity under Static Natural Scene Stimuli
In this experiment, we set all models to have 128-dimensional latent variables and train them for 5,000 iterations. The optimizer is set to Adam with a learning rate of 0.0001. At the training stage, each sample input to TiDeSPL-VAE is sequential neural activity from 5 time points. For self-supervised contrastive learning, the offset of positive samples from target samples is within $\pm$3 time points. At the inference stage, we consider the temporal dependence and the assumption that the latent variables of TiDeSPL-VAE are generated by an n-order Markov chain. Consequently, for a target time point, we form a sequence data including its antecedent n time points and its own to compute the latent variables (here n=4). The setup for the other models is given in Appendix E. To quantify performance in decoding natural scenes, we first obtain the latent variables of the last 20 time points (50ms-250ms) in each trial, since there is a response latency in the mouse visual cortex for static stimuli [36]. Then, these latent variables are concatenated into a vector as latent representations of neural activity for that trial. We use the KNN algorithm, a nonparametric supervised learning method, to classify the latent representations of each trial, i.e., to decode the corresponding natural scenes. We search the number of neighbors in odd numbers from 1 to 20 and use the Euclidean distance metric. We fit the KNN using the training set and choose the best number of neighbors on the validation set with classification accuracy as the metric. The accuracy of the test set is reported as the decoding score.
@@
Table 2: The decoding scores (%) for natural scene classification on the visual neural dataset. The standard error is computed based on 10 runs with different random initialization.
Models Mouse 1 Mouse 2 Mouse 3 Mouse 4 Mouse 5
PCA (baseline) 20.0 43.2 51.2 43.6 20.0
$\beta$-VAE 57.6$\pm$4.3 47.6$\pm$1.4 33.6$\pm$1.8 46.4$\pm$2.3 44.4$\pm$2.2
LFADS 55.2$\pm$2.8 50.0$\pm$3.1 48.8$\pm$2.5 54.8$\pm$1.8 50.4$\pm$2.5
pi-VAE 76.4$\pm$3.4 67.2$\pm$2.5 68.8$\pm$2.3 81.2$\pm$2.5 30.4$\pm$2.2
Swap-VAE 86.0$\pm$1.6 70.8$\pm$2.0 54.4$\pm$1.5 67.2$\pm$2.9 59.6$\pm$3.5
CEBRA 47.6$\pm$2.3 46.0$\pm$1.6 42.8$\pm$2.0 52.4$\pm$1.7 45.2$\pm$1.5
TiDeSPL-VAE 96.4$\pm$1.1 74.8$\pm$2.0 74.8$\pm$1.7 78.8$\pm$2.9 67.6$\pm$2.1
Figure 3: Results on the visual neural dataset under natural scenes (Mouse 1). All dimension reduction is done by tSNE. A. The upper panel is the 2-dimensional embedding of the latent representations from TiDeSPL-VAE for each trial. Each color corresponds to one natural scene. Transparent dots denote trials from the training set. Hollow dots denote trials from the test set. The lower panel is the 2-dimensional embedding of the latent trajectories over time, averaged over all trials for each scene. The asterisk is the starting point. B-E. The same visualization as A for alternative models.
Results of decoding scores
As shown in Table 2, TiDeSPL-VAE achieves the highest decoding scores on four mice with a noticeable improvement over other models. We observe some meaningful phenomena in the comparisons. The fact that pi-VAE introduces class labels of natural scenes in the training stage leads to high decoding scores, but not stably (Mouse 5). The performance of Swap-VAE is consistently good, which suggests that the swapping operation for training is indeed effective. However, the other two models (LFADS and CEBRA) that take sequential data as input instead perform poorly in the downstream task, suggesting that their approaches (bidirectional RNNs and temporal filters) for extracting temporal neural features are less suitable in this case. In contrast, the chronological stepwise computation in our model reliably captures time dependence.
Results of latent structures
In addition to the quantitative analysis, we visualize the latent representations by embedding them in two dimensions using tSNE (Figure 3, we focus on Mouse 1 with the highest scores. See Appendix G for results of other mice). On the one hand, we apply dimension reduction to the representations of each trial (including all time points from 50ms to 250ms), to show the trial-to-trial performance in decoding natural scenes. On the other hand, we reduce the dimensions of a single time point and take the average across trials, to show the latent trajectories over time. For the results of TiDeSPL-VAE, we first observe that most trials are well separated for different classes, especially those from the test set. The latent trajectories capture a similar clear temporal structure with weak class information. For pi-VAE and Swap-VAE, while they are able to discriminate scene classes of trials, the latter part of latent trajectories show varying degrees of entanglement across time. As for LFADS and CEBRA, their ability to encode temporal features of sequential neural activity results in explicit temporal evolution, but the latent representations of different classes are largely intermingled. These results suggest that our model effectively distinguishes between category information and explicitly captures temporal information from neural dynamics.
In summary, these results demonstrate the advantages of our model in decoding natural scene stimuli and extracting fine neural dynamics. However, we also see that the decoding performance for different mice exhibits large differences and the latent representations of some trials are not able to unravel the category entanglement. This may be attributed to the substantial variability in the neural activity of different mice and different trials during passive viewing.
Table 3: The decoding scores (%, in 1s window) for natural movie frame classification on the visual neural dataset. The standard error is computed based on 10 runs with different random initialization.
Models Mouse 1 Mouse 2 Mouse 3 Mouse 4 Mouse 5
PCA (baseline) 8.44 28.77 25.42 21.56 11.69
$\beta$-VAE 7.44$\pm$0.24 15.13$\pm$0.35 14.00$\pm$0.37 17.11$\pm$0.50 9.28$\pm$0.31
LFADS 8.94$\pm$0.25 26.57$\pm$2.46 26.77$\pm$2.23 24.76$\pm$1.80 12.69$\pm$1.38
pi-VAE 10.24$\pm$0.31 42.51$\pm$0.65 36.96$\pm$0.60 38.31$\pm$0.52 18.08$\pm$0.59
Swap-VAE 11.09$\pm$0.25 44.99$\pm$0.76 36.37$\pm$1.53 37.68$\pm$1.14 19.14$\pm$0.63
CEBRA 10.62$\pm$0.18 52.76$\pm$0.89 61.01$\pm$0.76 42.11$\pm$0.73 22.33$\pm$0.31
TiDeSPL-VAE 13.88$\pm$0.19 65.38$\pm$0.36 59.88$\pm$0.72 54.33$\pm$0.54 30.18$\pm$0.40
@@
Figure 4: Results on the visual neural dataset under natural movie (Mouse 2). All dimension reduction is done by tSNE. A. The 2-dimensional embedding of the latent representations from TiDeSPL-VAE for each frame in the middle 10s of the entire movie. Each color corresponds to all frames within 1s. Small dots denote one frame. Large dots denote the average among all frames within the 1s. The red dashed line connects all averages. B-C. The same visualization as A for alternative models. D. The decoding scores (%) for natural movie frame classification across different time window constraints. Error bars indicate the standard error for 10 runs.
4.3.2 Experiments for Neural Activity under Natural Movie Stimuli
Experiment setup and evaluation
In this experiment, all models are also set up with 128-dimensional latent variables and trained for 20,000 iterations. The optimizer and the learning rate are the same as for the dataset under natural scenes. For TiDeSPL-VAE, at the training stage, we use neural activity from 4 time points as one sample and set the maximum absolute offset to 2. At the inference stage, following the approach in Section 4.3.1, the latent variables of a target time point are generated based on antecedent n time points (n=3) and its own. The setup for the other models is also presented in Appendix E. To quantify the performance in decoding natural movie frames, we compute the latent variables of 4 time points within each frame and take the average as latent representations for that frame. KNN is applied to predict movie frames corresponding to latent representations (900 frames in total, i.e., 900 classes). Following a previous method [34], we take the accuracy measured by considering the error between a predicted frame and the true frame within 1s (default size of time window constraint) as a correct prediction. We similarly use the validation set to find the best number of neighbors and report the accuracy of the test set as the decoding score.
Results of decoding scores
We find that TiDeSPL-VAE performs best on four mice (Table 3). In particular, our model achieves significantly higher decoding scores than LFADS. Furthermore, we analyze the decoding scores under different sizes of time window constraints (maximum difference between predicted and real frames) on Mouse 2, since most models achieve the highest scores on it. Figure 4D shows that our model consistently outperforms CEBRA and Swap-VAE across a wide range of window sizes, while CEBRA’s performance degrades faster at smaller size constraints.
Results of latent structures
We similarly reduce the dimensions of latent representations of each movie frame for visualization. We set all frames within 1s as a group and show the trajectories of latent representations evolving over time for the middle 10s of the movie (Figure 4A-C, see Appendix G for results from other parts of the movie). Compared to the other models, the representations of TiDeSPL-VAE show clear temporal structure along movie frames and less overlap and entanglement between different groups.
To conclude, our model constructs meaningful latent representations related to the content and temporal structure of movie stimuli at large time scales. At small time scales, although there is a drop in decoding performance, our model still significantly outperforms the alternatives. Similar features of adjacent frames may be a factor that makes the latter problem challenging, which may require further exploration.
Table 4: The decoding scores (%) of ablation studies on the loss function and the recurrent module of TiDeSPL-VAE. The standard error is computed based on 10 runs.
Natural Scenes Natural Movie
Models Mouse 1 Mouse 2 Mouse 1 Mouse 2
TiDeSPL-VAE 96.4$\pm$1.1 74.8$\pm$2.0 13.88$\pm$0.19 65.38$\pm$0.36
Without negative samples 94.0$\pm$1.7 70.4$\pm$0.8 11.27$\pm$0.36 49.59$\pm$1.18
Without contrastive loss 89.2$\pm$1.5 68.8$\pm$1.9 11.24$\pm$0.23 47.98$\pm$0.90
Without swap operation 90.4$\pm$1.7 68.0$\pm$2.7 10.22$\pm$0.31 49.39$\pm$0.57
Without swap operation and constrastive loss 84.4$\pm$2.6 58.0$\pm$2.1 9.16$\pm$0.37 24.84$\pm$1.10
With temporal independent prior 87.2$\pm$1.9 70.8$\pm$3.3 12.09$\pm$0.16 57.74$\pm$0.62
[…]
4.4 Ablation Studies
We perform ablation studies in several aspects of TiDeSPL-VAE’s components and neural activity input dimensions to explore their impact on performance. We present the results of Mouse 1 and Mouse 2 in the main text and the results of the other mice are given in Appendix H.
The loss function and the recurrent module of TiDeSPL-VAE
To show the effectiveness of components of our model, we conduct some ablation studies on the loss function and the recurrent module (Table 4 and 7). In terms of contrastive learning, we first exclude negative samples from the computation of the contrastive loss and use only the cosine distance between the content latent variables of the positive sample pairs as the loss function, i.e., only bring the positive pairs closer. There is a slight decrease in model performance, suggesting that negative samples are useful. Then we directly remove the constrastive loss or the swap operation, both of which show a similar impact. We last remove both, i.e., there are no more objectives related to constrastive learning. The large drop in performance suggests that contrast learning plays a crucial role in our models. In terms of the regular loss of style latent variables, we originally assumed that the prior distribution is time-dependent. When we assume that it is an independent standard normal distribution at each time step, the model performance degrades, demonstrating that time-dependent assumptions about the prior are also important. In terms of the recurrent module, the results suggest that GRU is a better choice in considering the trade-off between performance and computational efficiency.
Table 5: The decoding scores (%) of ablation studies on the content and style latent representations.
Natural Scenes Natural Movie
Mouse 1 Mouse 2 Mouse 1 Mouse 2
Content 94.0$\pm$1.3 70.8$\pm$2.1 14.02$\pm$0.27 68.77$\pm$0.46
Style 76.4$\pm$1.9 62.0$\pm$2.5 7.48$\pm$0.20 14.87$\pm$0.73
Figure 5: The results of ablation studies on the dimension of latent variables and the number of input neurons. Error bars indicate the standard error for 10 runs.
The content and style latent representations
We evaluate decoding scores of content and style latent representations (Table 5 and 8). We find that content variables stably outperform style variables, supporting our definition of the two, i.e., content variables are more relevant to visual stimuli.
The dimension of latent variables and the number of input neurons
We perform ablation studies on the number of latent variables and input neurons. As shown in Figure 5 and 9, first, the model performance saturates gradually as the dimension of latent variables increases, which suggests that it is sufficient to choose a dimension at reasonable intervals (not too much fewer than input neurons). Second, performance decreases as the number of sampled neurons decreases, suggesting that for each mouse, all recorded neurons contribute to the representation of visual stimuli.